OpenStudy (anonymous):

I cannot find these things anywhere in my book, for even an example... Over two summers, Ray saved $800.00 and $600.00. The polynomial represents his savings at the beginning of the third year, where x is the growth factor. (The interest rate r is x – 1.) What is the interest rate he needs to save $1,650.00 after the third summer? 800x2 + 600x (1 point) 0.1% 10.9% –285.9% 1.1%

OpenStudy (anonymous):

what are those number after (1 point) rewrite the question. if those are potential answers number them or denote them somehow.

OpenStudy (phi):

this is a goofy way of asking to solve for x in the quadratic \[ 800 x^2 +600x = 1650 \] and then find the interest rate r = x – 1 Write r as a percent (i.e. 100*(x-1) will be the answer)

OpenStudy (anonymous):

ivan77 those are the possible answer. It's multiple choice.

OpenStudy (anonymous):

phi I kind of understand what you put down, but not quite.

OpenStudy (phi):

The polynomial represents his savings at the beginning of the third year those words mean that \[ 800x^2 + 600x\] is a way to calculate his savings at the start of the 3rd year x is how much the savings grow. They want you to find x so that the expression equals 1650

OpenStudy (anonymous):

So I have to get x by itself?

OpenStudy (phi):

yes, you have to solve a quadratic equation 800 x^2 +600x = 1650 you can divide by 10 (all terms) to simplify it a bit \[ 80 x^2 +60 x - 165 =0 \]

OpenStudy (anonymous):

I can't find an example of how to solve a quadratic equation in my book...

OpenStudy (phi):

what are you studying ?

OpenStudy (anonymous):

Algebra II

OpenStudy (phi):

Solving quadratics is a part of Algebra II Are you allowed a calculator?

OpenStudy (anonymous):

Imean I'm home schooled I'll do whatever I want. So yes I can use a calculator.

OpenStudy (phi):

Ok, take a look at the video. It should be helpful.